Project

General

Profile

Actions

20180220

Location : MICE meeting room R1 2nd Floor East at 14:00 pm

Phone access : phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

  • February jobs:
    • Tracker:
      • Preparations now being made for the survey of the trackers within the bore. Presently John, Craig and Kevin are preparing for the installation of the pieces that hold the prisms for alignment.
      • Cryostats have been electrically disconnected. Waveguides downstream are disconnected.
      • EMR support frame is still in place.
      • Cryostats will be moved today, Cryo-3 has been moved to a temporary location.
      • Filtered lights will be installed and then the survey points will be installed.
      • Survey of tracker scheduled for tomorrow.
      • Tony Millington will be the master of the tracker position.
      • Helium window measurement will follow once the FC has been extracted. Need to check withTim Hayler re drawings and acceptance metrology.
    • Diffuser:
      • Diffuser extraction (survey complete) -- extraction to be carried out once SSD is extracted.
      • Need to verify thickness and composition of the diffuser.
    • Target Controls test #1954 (Henry/Paolo 21st)
      • Will execute target controls tomorrow. Goal is to reboot target PC and target rack.
    • D2 field mapping (Paolo/Henry/??? 28th)
      • Repeat D2 field mapping; propose to do on 28th because this is the maintenance day. Health Physics has beam contacted to do the necessary radiation survey. Arrangements to be confirmed with the DO.
      • Also wish to measure the D1 field. Evidence that box is magnetic . Need to make a plan for how measurement is to be made. Deadline for plan is the next Ops meeting.
    • https://micewww.pp.rl.ac.uk/attachments/download/9660/2017-12-06_decommissioning-measurements_v3.pdf
      • Document stands; no updates.
  • March jobs:
    • SSU, SSD Magnet field map. 19th to 23rd March
      • CW needs to work out access passes for the CERN staff. Field map will be done at 2T (SSD, M2 on), then ECE to 3T. Again, need to make a plan to ensure best use of CERN magnet-team's time. SSU at 3T, and 4T for VB.
    • ISIS Target Training
      • ISIS nominated target expert will be available at the MICE/ISIS safety meeting on Thursday. Idea probably that operation of target will be in March during the Cycle.
  • Next meeting - 06Mar18
  • AOB

Updated by Nebrensky, Henry over 4 years ago · 8 revisions